# Generative modelling with jump-diffusions

### v1

@@ 3.1 The jump-diffusion Laplace model
$G_{2}(x,t)=\frac{\partial}{\partial x}\mathcal{F}^{-1}\left\{\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}\right\}$ (19)
Explicit integral expressions for $G_{1,2}(x,t)$ are given in Eqs. (64,66). Evaluating these integrals is computationally costly, but since $\mathbf{f},g$ are typically unchanged for a given class of models (independent of $p_{\rm data}$), it is advantageous to create numerical representations of these functions offline and use them like analytical functions in different applications. This can be done by calculating the integrals at discretized points in the range $t\in[0,T]$ and $x\in[0,x_{\rm max}]$ and defining function approximations by interpolation.
From Eqs. (16,17) one obtains the FT of the stationary distribution $p_{\rm st}(x)=\lim_{t\to\infty}p(\mathbf{x},t|\mathbf{x}_{0})$ as
$\hat{p}_{\rm st}(k)=\left(1+\frac{\sigma^{2}}{2}k^{2}\right)^{-\lambda}e^{-\frac{D}{2}k^{2}}$ (20)
While the inverse FT is also here not available in closed form, it is straightforward to sample from $p_{\rm st}(x)$, see Eq. (25).
The JDL model implementation thus requires the following steps:
  1. 1.
    Evaluating the functions $G_{1,2}(x,t)$ numerically in terms of the integrals given in Eqs. (64,66) over the range $t\in[0,T]$ and $x\in[0,x_{\rm max}]$.
  2. 2.
    Optimizing Eq. (12) with the numerical representations of $p(\mathbf{x},t|\mathbf{x}^{\prime})$ and $\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ to obtain $\mathbf{s}_{\boldsymbol{\theta}}$ as approximation of the generalized score function. The Monte-Carlo integration requires simulation of Eq. (15), which is simplified due to the linear form of the dynamics. Direct integration yields
    $\mathbf{Y}(t)=\mathbf{Y}(0)e^{-t/2}+\bar{\boldsymbol{\xi}}_{-}^{\rm G}(t)+\bar{\mathbf{J}}_{-}(t),$ (21)
    where I define
    $\bar{\boldsymbol{\xi}}_{\pm}^{\rm G}(t)=\int_{0}^{t}\mathrm{d}s\,e^{\pm(t-s)/2}\boldsymbol{\xi}^{\rm G}(s),\qquad\qquad\bar{\mathbf{J}}_{\pm}(t)=\sum_{j=1}^{N_{t}}\mathbf{A}_{j}e^{\pm(t-\tau_{j})/2}.$ (22)
    Eq. (21) allows for exact sampling without discretization error, see Point 4 below for details.
  3. 3.
    Implementing the generative process via the probability flow ODE Eq. (10) or SDE Eq. (11) upon replacing the generalized score function by $\mathbf{s}_{\boldsymbol{\theta}}$. Efficient sampling can be implemented with exponential samplers [21] that exploit the linearity akin to Eq. (21), only that now a time discretization is needed due to the nonlinear drift. I use the simple discretizations:
@@
    for Eq. (10), and
    $\mathbf{X}_{i+1}=\mathbf{X}_{i}e^{\Delta t/2}+4\left(e^{\Delta t/2}-1\right)\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i})+\bar{\boldsymbol{\xi}}_{+}^{\rm G}(\Delta t)+\bar{\mathbf{J}}_{+}(\Delta t)$ (24)
    for Eq. (11). In Eqs. (23,24), the starting point is sampled $\mathbf{X}_{0}=\mathbf{X}(0)\sim p_{st}(x)$.
  4. 4.
    Sampling implementations:
@@
[…]
      where $\mathbf{Z}_{1}\sim\mathcal{N}(\mathbf{0},D\mathbf{I})$, $\mathbf{Z}_{2}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$, and $\Gamma\sim\text{Gamma}(\lambda)$.
@@
      Gaussian noise increments $\bar{\boldsymbol{\xi}}_{\pm}^{\rm G}(t)$. The definitions Eq. (22) imply $\bar{\boldsymbol{\xi}}_{-}^{\rm G}(t)\sim\mathcal{N}(\mathbf{0},D\left(1-e^{-t}\right))$ and $\bar{\boldsymbol{\xi}}_{+}^{\rm G}(t)\sim\mathcal{N}(\mathbf{0},D\left(e^{t}-1\right))$.
@@
[…]
4 Connection with LIM
It is instructive to compare the results in this work with those of the LIM framework of [20], where the forward process Eq. (1) is driven by $\alpha$-stable Lévy noise with $21<\alpha<$ 2. In this case, a modification of the standard score function has also been obtained leading to the fractional score function $\frac{\Delta^{\frac{\alpha-2}{2}}\nabla p(\mathbf{x},t)}{p(\mathbf{x},t)}$, where $\Delta^{\frac{\alpha-2}{2}}$ denotes the fractional Laplacian of order $\frac{\alpha-2}{2}$. While the derivation of the jump-diffusion generative process in Appendix A is not valid for this case, one can naively consider $\alpha$-stable noise in the generalized score function by substituting the corresponding $\psi(k)$ in Eq. (9). Indeed, the characteristic function of isotropic $\alpha$-stable noise increments $\Delta\boldsymbol{\xi}$ with $21<\alpha<$ 2 is likewise given by Eq. (7), but with $\psi(k)=k^{\alpha}$ [4]. For this special case, one finds from Eq. (9)
@@ 4 Connection with LIM
and the generalized score function $\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}$ becomes the fractional score function derived in [20]. The generative process Eq. (10) then recovers the probability flow formulation of LIM [20]. These results are somewhat surprising, since the rigorous analysis of the the time-reversal of Markov jump processes in [3] shows that in the case of $\alpha$-stable Lévy processes with $21<\alpha<$ 2, a very different time-reversal formula applies than for finite activity Lévy processes. However, the results for the generative process are consistent, which suggests that the generalized score function with $\mathcal{V}(\mathbf{x},t)$ of Eq. (9) might be valid more generally.
Specifying the forward process further as an Ornstein-Uhlenbeck process with $\mathbf{f}(\mathbf{x},t)=\frac{1}{2}\mathbf{x}$ and $g(t)=1$, it is also straightforward to show that the conditional generalized score function for $\psi(k)=k^{\alpha}$ has the form (see Appendix C.2)
@@
$\frac{\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})}{p(\mathbf{x},t|\mathbf{x}^{\prime})}$ $=$ $-\frac{\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}{2\left(1-e^{-t}\right)}$ (27)
recovering both the familiar Gaussian result for $\alpha=2$ and the result for the conditional fractional score function derived in [20] for $21<\alpha<$ 2 using a more cumbersome method. All power-law forms $\psi(k)\propto k^{\alpha}$ would thus lead to a simple functional form of the conditional generalized score function, if the validity of Eq. (9) could be extended to these cases.
5 Numerical experiment: heavy-tailed data distribution
(a)           (b) Method MSLE SE JDL-SDE 0.018 6.52×$\times 10^{-4}$ JDL-ODE 0.040 3.68×$\times 10^{-3}$ DLPM ($\alpha=1.7$) [14] 0.071 0.028 LIM ($\alpha=1.7$) [14] 0.267 0.077 DDPM [14] 0.528 0.4
Table 1: (a) Heavy-tailed samples from $p_{\rm data}$ given as an isotropic $\alpha$-stable distribution in $d=2$ as $\mathcal{F}^{-1}\left\{e^{-k^{\alpha}}\right\}$ with $\alpha=1.7$. Only one component is shown. (b) Measured values for the MSLE Eq. (28) over 20 runs. JDL-SDE and JDL-ODE values are calculated from 25,000 samples generated by Eqs. (23,24). The other values are taken from [14], whose $\alpha$-stable extension of DDPM is referred to as denoising Lévy probabilistic model (DLPM). For more details on the implementation, see Appendix F.
In order to demonstrate the applicability of the framework, I consider a simple toy experiment following [14], in which $p_{\rm data}$ is given by an $\alpha$-stable distribution with $\alpha=1.7$ in $d=2$. With such a strongly non-Gaussian data distribution one can test how well the generative process is able to capture heavy tails in the distribution, which is challenging with Gaussian methods. A suitable metric to compare the performance of different methods in approximating the tails is the mean-square logarithmic error (MSLE) of the higher quartiles, defined as [14]
${\rm MSLE}(\xi)=\int_{\xi}^{1}\mathrm{d}p\left(\log F_{\rm data}^{-1}(p)-\log F_{\rm gen}^{-1}(p)\right)^{2}$ (28)
where $\xi$ has been chosen as 0.95 and $F^{-1}$ denotes the inverse cumulative distribution function. For the implementation of JDL, Eq. (28) is evaluated separately for the two components and then averaged. As the results in Table 1 show, the JDL implementations in ODE and SDE form, Eqs. (23,24) strongly outperform both LIM [20] and DDPM with $\alpha$-stable noise [14]. In addition, the standard error of the samples is considerably smaller, by one magnitude for JDL-ODE and two magnitudes for JDL-SDE. It should be noted that neither the parameters of the jump-diffusion process (jump rate $\lambda$ and variance $\sigma^{2}$) nor the neural network used to estimate the generalized score function have been optimized in any way. Nevertheless, the JDL model provides better tail approximation than the two heavy-tailed methods with parameter $\alpha=1.7$, that is, set to the same stability index as the data distribution, where optimal performance might be expected. For more details on the implementation, see Appendix F

### v2

@@ 3.1 The JL model: Laplace distributed jump amplitudes
$\rho_{\rm L}(\mathbf{x},\sigma)=\frac{2}{(\sigma)^{d}(2\pi)^{d/2}}\left(\frac{\|\mathbf{x}\|}{\sqrt{2}\,\sigma}\right)^{1-\frac{d}{2}}K_{1-\frac{d}{2}}\left(\frac{\|\mathbf{x}\|\sqrt{2}}{\sigma}\right)$ (18)
is the distribution $\text{L}_{d}(\sigma^{2})$ itself (see Appendix D). The conditional generalized score function is
$\mathcal{S}(\mathbf{x},t|\mathbf{x}^{\prime})=\frac{\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}}{\|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\|}\mathcal{G}(\|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\|,t)$ (19)
introducing the function
$\mathcal{G}(x,t)=-\frac{1}{2}x-\frac{\sigma}{\sqrt{2}}\frac{K_{d/2}\left(\frac{x\sqrt{2}}{\sigma}\right)}{K_{1-d/2}\left(\frac{x\sqrt{2}}{\sigma}\right)}\left(\frac{e^{-t}}{1-e^{-t}}\right),\qquad\qquad x>0$ (20)
with $\mathcal{G}(x,t)=0$ when $x=0$. In Eqs. (18,20), $K_{\nu}$ denotes the modified Bessel function of the 2nd kind. Moreover, the stationary distribution is just
$p_{\rm st}(\mathbf{x})=\lim_{t\to\infty}p(\mathbf{x},t|\mathbf{x}^{\prime})=\rho_{\rm L}(\mathbf{x},\sigma)$ (21)
and can be sampled like the jumps.
3.2 Implementation of the JL model
The implementation of the JL model thus requires the following steps:
1. Estimating the generalized score function. Optimizing the loss function Eq. (13) with $\mathcal{S}(\mathbf{x},t|\mathbf{x}^{\prime})$ given in Eq. (19) yields $\mathbf{s}_{\boldsymbol{\theta}}$ as approximation of the exact $\mathcal{S}(\mathbf{x},t)$. As in the Gaussian case, the numerical stability of the training process can be improved by scaling the time-dependence in the score function. Eq. (20) indicates that the divergence in $\mathcal{S}(\mathbf{x},t|\mathbf{x}^{\prime})$ for $t\to 0$ can be removed by multiplying with $1-e^{-t}$, thus the loss function becomes
$J_{\rm JD}(\boldsymbol{\theta})=\int_{0}^{T}\mathrm{d}t\frac{w(t)}{(1-e^{-t})^{2}}\mathbb{E}_{\mathbf{Y}_{0}\sim p_{\rm data}}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\left\|\hat{\mathbf{s}}_{\boldsymbol{\theta}}(\mathbf{Y},t)-(1-e^{-t})\mathcal{S}(\mathbf{Y},t|\mathbf{Y}_{0})\right\|^{2}\right]$ (22)
Direct integration of Eqs. (1,6) with $\mathbf{f}(\mathbf{x},t)=-\frac{1}{2}\mathbf{x}$, $g(t)=1$, $D=0$ yields
$\mathbf{Y}(t)=\mathbf{Y}(0)e^{-t/2}+\mathbf{J}(t)$ (23)
where the jump increment is derived as
$\mathbf{J}(t)=\sum_{j=1}^{N_{t}}\mathbf{A}_{j}e^{-(t-\tau_{j})/2}$ (24)
Eqs. (23,24) allow for exact sampling of $\mathbf{Y}(t)$ without discretization errors, see Section 3.3 below for details. Setting the weight function $w(t)=(1-e^{-t})^{2}/T$ in Eq. (22) then yields the loss function
$J_{\rm JD}(\boldsymbol{\theta})=\mathbb{E}_{t\sim{\rm U}(0,T)}\mathbb{E}_{\mathbf{Y}_{0}\sim p_{\rm data}}\mathbb{E}_{\mathbf{J}(t)}\left[\left\|\hat{\mathbf{s}}_{\boldsymbol{\theta}}(\mathbf{Y}_{0}e^{-t/2}+\mathbf{J}(t),t)-\frac{\mathbf{J}(t)}{\|\mathbf{J}(t)\|}\hat{\mathcal{G}}(\|\mathbf{J}(t)\|,t)\right\|^{2}\right]$ (25)
using Eqs. (19,23) and defining $\hat{\mathcal{G}}(x,t)=(1-e^{-t})\mathcal{G}(x,t)$.
2. Sampling of the generative process. Implementing the generative process upon replacing $\mathcal{S}$ by $\mathbf{s}_{\boldsymbol{\theta}}$ via the probability flow ODE Eq. (11) or SDE Eq. (12) can be made more efficient with exponential samplers [28] that exploit the linearity akin to Eq. (23). For Eq. (11), direct integration of the linear term while keeping $\mathbf{s}_{\boldsymbol{\theta}}$ constant yields the simple discretization:
@@ 3.2 Implementation of the JL model
$=$ $\mathbf{X}_{i}e^{\Delta t/2}+2\left(\frac{e^{\Delta t/2}-1}{1-e^{-(T-t_{i})}}\right)\hat{\mathbf{s}}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i})$
using the rescaled estimated score function introduced in Eq. (22). For Eq. (12) direct integration yields likewise
$\mathbf{X}_{i+1}=\mathbf{X}_{i}e^{\Delta t/2}+4\left(\frac{e^{\Delta t/2}-1}{1-e^{-(T-t_{i})}}\right)\hat{\mathbf{s}}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i}))+\tilde{\mathbf{J}}(\Delta t),$ (27)
where the (backwards) jump increment $\tilde{\mathbf{J}}(t)$ is derived as
$\tilde{\mathbf{J}}(t)=\sum_{j=1}^{N_{t}}\mathbf{A}_{j}e^{(t-\tau_{j})/2}$ (28)
In both Eqs. (26,27), the starting point is sampled $\mathbf{X}_{0}\sim\text{L}_{d}(\sigma^{2})$ since $p_{\rm st}(\mathbf{x})=\rho_{\rm L}(\mathbf{x},\sigma)$.
3.3 Sampling algorithms
The JL model requires the sampling of three different random variates:
@@ 3.3 Sampling algorithms
    $\mathbf{X}_{0}\sim\text{L}_{d}(\sigma^{2})$. As outlined in Appendix D, random variates of $\text{L}_{d}(\sigma^{2})$ are generated by setting
    $\mathbf{X}_{0}=\sqrt{\Gamma}\,\mathbf{Z}$ (29)
    where $\Gamma\sim\text{Gamma}(1)$ and $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$.
@@
    Jump increment $\mathbf{J}(t)$. The definition Eq. (24) can be implemented directly by: (i) Drawing $M\sim\text{Poisson}(\lambda t)$, where $\lambda=1$. (ii) Determining $M$ times $\tau_{j}\sim\text{U}(0,t)$ and $\mathbf{A}_{j}\sim\text{L}_{d}(\sigma^{2})$. (iii) Setting $\mathbf{J}(t)=\sum_{j=1}^{M}\mathbf{A}_{j}e^{-(t-\tau_{j})/2}$. However, a simpler method is found by noting that the distribution of $\mathbf{J}(t)$ is simply (compare Eqs. (17,23))
    $p(\mathbf{J},t)=\rho_{\rm L}(\mathbf{J},\sigma)\left(1-e^{-t}\right)+\delta(\mathbf{J})e^{-t}$ (30)
    Thus we can sample $\mathbf{J}(t)$ by selecting $\omega\sim U(0,1)$ and setting $\mathbf{J}\sim\text{L}_{d}(\sigma^{2})$ if $\omega>e^{-t}$ and $\mathbf{J}=\mathbf{0}$ otherwise.
@@
    Jump increment $\tilde{\mathbf{J}}(t)$. The definition Eq. (28) indicates that the same algorithm as in (b) applies upon changing (iii) to $\tilde{\mathbf{J}}(t)=\sum_{j=1}^{M}\mathbf{A}_{j}e^{(t-\tau_{j})/2}$. Interestingly, a simplified algorithm can here also be found. As derived in Appendix E, the PDF of $\tilde{\mathbf{J}}(t)$ is given by
    $p(\tilde{\mathbf{J}},t)=\rho_{\rm L}(\tilde{\mathbf{J}},\sigma\,e^{t/2})\left(1-e^{-t}\right)+\delta(\tilde{\mathbf{J}})e^{-t}$ (31)
    The random variates $\tilde{\mathbf{J}}(t)$ can thus be sampled by selecting $\omega\sim U(0,1)$ and setting $\tilde{\mathbf{J}}\sim\text{L}_{d}(\sigma^{2}e^{t})$ if $\omega>e^{-t}$ and $\tilde{\mathbf{J}}=\mathbf{0}$ otherwise.
4 Connection with other works
LIM. It is instructive to compare the results in this work with those of LIM [27], where the forward process Eq. (1) is driven by $\alpha$-stable Lévy noise with $1<\alpha\leq 2$. In this case, a modification of the standard score function has also been obtained leading to the fractional score function [27]
$\mathcal{S}_{\alpha}(\mathbf{x},t)=\frac{\Delta^{\frac{\alpha-2}{2}}\nabla p(\mathbf{x},t)}{p(\mathbf{x},t)}$ (32)
where $\Delta^{\frac{\alpha-2}{2}}$ denotes the fractional Laplacian of order $\frac{\alpha-2}{2}$. While the derivation of the jump-diffusion generative process in Appendix A is not valid for this case, one can naively consider $\alpha$-stable noise in the generalized score function by substituting the corresponding $\psi(k)$ in Eq. (10). Indeed, the characteristic function of isotropic $\alpha$-stable noise increments $\Delta\boldsymbol{\xi}$ with $1<\alpha\leq 2$ is likewise given by Eq. (7), but with $\psi(k)=k^{\alpha}$ [6]. For this special case, one finds from Eq. (10)
@@ 4 Connection with other works
and the generalized score function $\mathcal{S}(\mathbf{x},t)$ of Eq. (9) recovers the fractional score function Eq. (32). The generative process Eq. (11) then yields a probability flow ODE for LIM. These results are somewhat surprising, since the rigorous analysis of the time-reversal of Markov jump processes in [5] shows that in the case of $\alpha$-stable Lévy processes with $1<\alpha\leq 2$, a very different time-reversal formula applies than for finite activity Lévy processes. However, the result in Eq. (33) suggests that the generalized score function Eq. (9) might be valid more generally.
Generator matching. The generator matching framework of [9] specifies the backward process in terms of the generator of a general Markov process and learns its coefficients using a prescribed conditional probability distribution that maps between the simple distribution and $p_{\rm data}$ (called probability path). Like in flow-matching, the probability path replaces the forward noising process and is explicitly specified at model construction. The key insight in generator matching is that the probability path and conditional generator are related through the Kolmogorov-Forward equation. Thus the coefficients in the conditional generator can be determined once a path is given, although this is strictly only possible in $d=1$, while for $d>1$ a composition rule needs to be applied [9].
  1. 1.
    The framework here is consistent with generator matching if we take as probability path the time reverse of the conditional PDF $p(\mathbf{x},t|\mathbf{x}^{\prime})$ of the forward process Eqs. (1,6) and consider a backward generator containing a flow component only. The conditional flow field $\mathbf{u}(\mathbf{x},t|\mathbf{x}^{\prime})$ then satisfies the continuity equation (i.e., the adjoint Kolmogorov-Forward equation in this case) with $\tilde{p}(\mathbf{x},t|\mathbf{x}^{\prime})=p(\mathbf{x},T-t|\mathbf{x}^{\prime})$:
@@
    $\frac{\partial}{\partial t}\tilde{p}(\mathbf{x},t|\mathbf{x}^{\prime})=-\nabla\left[\mathbf{u}(\mathbf{x},t|\mathbf{x}^{\prime})\tilde{p}(\mathbf{x},t|\mathbf{x}^{\prime})\right]$ (34)
    Since $\tilde{p}(\mathbf{x},t|\mathbf{x}^{\prime})$ satisfies the same Fokker-Planck equation as the marginal PDF, the derivation in Appendix A can be performed in the same way and yields
    $\mathbf{u}(\mathbf{x},t|\mathbf{x}^{\prime})=-\mathbf{f}(\mathbf{x})+\mathcal{S}(\mathbf{x},T-t|\mathbf{x}^{\prime})$ (35)
    where the conditional generalized score function is defined via Eq. (14). Eq. (35) recovers the ODE representation Eq. (19). We see that generator matching reverses the procedure used here and yields the same results in arbitrary $d$ because the probability path is the solution of a given Fokker-Planck equation and thus the analytical results derived here can be used.
  2. 2.
    The SDE representation then likewise follows due to the fact that we can define a generator $\mathcal{L}_{t}^{\rm NG}$ satisfying
    $\mathcal{L}_{t}^{\rm NG}\tilde{p}(\mathbf{x},t)$ $=$ $-\nabla\cdot\left[\mathcal{S}(\mathbf{x},T-t)\tilde{p}(\mathbf{x},t)\right]+\frac{D}{2}g^{2}(t)\nabla^{2}\tilde{p}(\mathbf{x},t)$ (36)
    $+\lambda\int\mathrm{d}\mathbf{z}\,\rho(z)\left[\tilde{p}(\mathbf{x}+g(t)\mathbf{z},t)-\tilde{p}(\mathbf{x},t)\right]$
    $=$ $0$ (37)
    where $\mathcal{S}(\mathbf{x},t)$ is defined in Eq. (9) and $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$. Eq. (37) likewise follows from the results in Appendix A. In generator matching, a generator satisfying $\mathcal{L}_{t}\tilde{p}(\mathbf{x},t)=0$ is called a divergence-less generator, which can be added to a pure flow ODE model to turn it into a model with stochastic sampling as in the SDE representation Eq. (12). Crucially, since the flow component is the same as for the ODE, this does not require estimation of additional coefficients. The generator Eq. (36) can be combined with other generators according to the generator matching composition rules to formulate more complex models.
  3. 3.
    It should be highlighted that, without a divergence-less generator, generator matching with jumps always contains an estimated jump distribution for the backward process, since the jumps need to be adapted to the chosen probability path. As shown in [9], even simple probability paths then lead to complicated jump distributions with space- and/or time-inhomogeneities.
5 Numerical experiments
Figure 1: Numerical experiments in $d=2$: (a,b,c) $p_{\rm data}$ given as a standardized swiss roll distribution; (d,e,f) $p_{\rm data}$ given as a Gaussian mixture. Plotted is the $F_{1}$ metric that measures the accuracy of the generated distributions. While the best performance is achieved by the Gauss-ODE method with $\max F_{1}>80$ in both tests, the JL model performs better at high noise intensities and for small step numbers. JL performs consistently better than LIM-SDE apart from the regime of very low noise intensities. Protocol details: $\alpha=1.9$ for LIM-SDE, $10^{5}$ training samples, $2\cdot 10^{4}$ test samples, $T=10$, error bars given as the standard error from 5 repetitions, sampling steps =100=100 in (a,d). The neural network is a simple multilayer perceptron with 4 layers, 128 nodes, and GELU activations. For more details on the implementation and the Gauss-ODE and LIM-SDE methods, see Appendix F.
In order to demonstrate the applicability of the framework, I consider numerical experiments on two simple data distributions in $d=2$: (a) a standardized swiss roll distribution; and (b) a mixture of 9 Gaussians with varying weights arranged on a grid. The focus is here to understand the effect of non-Gaussian jumps on the accuracy of sample generation compared with an equivalent Gaussian DM. Crucially, the underlying forward process, training protocol, and neural network architecture are intentionally kept very simple to disentangle effects due to specific implementations and those due to the noise type only. Indeed, it is likely that non-Gaussian DMs require different types of protocols/architectures to perform optimally compared with standard DMs, whose designs have been optimized for various applications in recent years.
Fig. 1 shows results for the JL model (both ODE and SDE formulations), the corresponding Gaussian DM as a probability flow ODE, and an equivalent LIM SDE of [27], which is driven by isotropic $\alpha$-stable noise. More details on the numerical implementations are presented in Appendix F. As performance metric I choose the $F_{1}$ metric, which is based on an extension of the notions of precision and recall to continuous distributions [19], and Fig. 1 shows the maximal $F_{1}$ values for samples generated by the four different methods.
In Figs. 1(a,d) results for fixed number of sampling steps (=100=100) and different noise intensities $D,\sigma^{2}=0.001-10$ are shown. For LIM-SDE the noise intensity is undefined, but scaling the noise term by $(\sigma^{2}/2)^{1/\alpha}$ ensures that for $\alpha=2$, i.e., when the $\alpha$-stable noise converges to Gaussian white noise, the resulting noise intensity is $\sigma^{2}$. The non-Gaussian methods all exhibit non-monotonic behaviour, where optimal performance is achieved in the range $\sigma^{2}=0.01-0.1$, but still below that of the Gaussian model at same noise intensity. For larger noise intensity the accuracy of all methods drops considerably. Interestingly, we can identify a parameter regime, where JL-ODE performs significantly better than Gauss-ODE, which is at $D,\sigma^{2}=1$ (Fig. 1(c,f)) and small number of sampling steps. For steps=25=25, JL-ODE achieves $\max F_{1}$ of 0.66 for the swiss roll and 0.55 for the Gaussian mixture, while $\max F_{1}<0.52$ and <0.37<0.37, respectively, for the other methods. For $D,\sigma^{2}<1$ on the other hand, Gauss-ODE performs best throughout, achieving the highest absolute values of $\max F_{1}\geq 0.8$ in both tests. Comparing the JL models and LIM-SDE, we see that LIM-SDE performs better at very low noise intensities and also achieves higher absolute $\max F_{1}$ values, but is consistently outperformed at larger noise intensities. In particular, for $D,\sigma^{2}=1$, LIM-SDE performs quite poorly compared with the JL models. It should be noted that the same Gaussian mixture test of Fig. 1(d,e,f) was first implemented in [20], where LIM-SDE was also tested and achieved $\max\,F_{1}=0.88$ for $\alpha=1.9$, but for a more sophisticated neural network architecture. The DLPM method of [20] achieved $\max\,F_{1}=0.94$, but its performance under the test conditions here was not evaluated.